SME Company has a debt-equity ratio of .60. Return on assets is 7.9 percent, and total equity is $510,000. What is the equity multiplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) What is the return on equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16) What is the net income? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) a. b. c. a. Equity multiplier b. Return on equity c. Net income times
Answer #1

1.Equi Multiplier = 1+ debt/Equity. Since debt-equity is 0.6, The equity multiplier is 1+0.6 = 1.60

2. Debt/Equity = 0.6 and Equity = 510,000. So debt is 0.6*510,000 = 306,000. So Total debt + equity = 306,000 + 510,000 = 816,000.

So total assets = Total debt and equity = 816,000

hence net profit = 816,000*0.079 = 64,464

Therefore ROE = 64,464/510000 = 0.1264= 12.64%

3. The net Profit = net income = $64,464.

Thanks

Let me know if you have any doubt